OpenStudy (owlstudies):

Complete the Square:

OpenStudy (owlstudies):

\[2x^2+8x-7=-2\]

OpenStudy (yowarning):

break it down to a simply problem

OpenStudy (yowarning):

then do the pemdas stage

OpenStudy (owlstudies):

would you start by adding seven to both sides?

OpenStudy (mathmale):

\[2x^2+8x-7=-2\]

OpenStudy (mathmale):

You could do that, but I don't see how that step would help you. This is an equation, namely, a quadratic equation, and you'll want all terms on the left side.

OpenStudy (mathmale):

If you want to move the -2, fine. Add 2 to both sides of this equation.

OpenStudy (owlstudies):

so 2x^2+8x-5=0

OpenStudy (mathmale):

Yes. Now make things easier for yourself by factoring 2 out of the first 2 terms. Use parentheses. Leave the -5 alone.

OpenStudy (owlstudies):

2(x^2+4x)-5=0?

OpenStudy (mathmale):

Yes, very good. Now focus on completing the square of x^2+4x.

OpenStudy (mathmale):

1) Take half of the coefficient of x (which is 4). 2) Square your result. 3) Add this result to the first 2 terms, and then subtract this result from the left 3 terms. Show your work.

OpenStudy (owlstudies):

I'm confused. what terms do i have to add or subtract it from? i get the whole (b/2)^2 , I'm not sure on where to add this too

OpenStudy (mathmale):

b is 4. Half of b is 2. Square this result. It's 4. Add this 4, and then subtract this 4: x^2+4x +4 -4

OpenStudy (mathmale):

x^2+4x+4 is the square of what binomial? Write your answer in the form ( )^2-4.

OpenStudy (owlstudies):

(x+2)^2-4

OpenStudy (mathmale):

Yes. Now, let's put this problem back together: 2x^2+8x-5=0 2(x^2+4x)-5 = 0 Replace (x^2+4x) with ( (x+2)^2 -4 ). Show your work.

OpenStudy (owlstudies):

so it'd be 2(x+2)^2-9=0?

OpenStudy (mathmale):

Here's what I would do: 2 [(x^2+4x+4 - 4)] - 5 = 0 2 [ (x+2)^2 - 4] - 5 = 0 2(x+2)^2 - 8 -5 = 0 Please go thru this and decide whether you can accept it or not.

OpenStudy (owlstudies):

okay i think i got it. thank you

OpenStudy (mathmale):

My pleasure! Bye for now.